# From Individual Experience to Collective Evidence: A Reporting-Based Framework for Identifying Systemic Harms

### v1

@@ Abstract
When an individual reports a negative interaction with some system, how can their personal experience be contextualized within broader patterns of system behavior? We study the incident database problem, where individual reports of adverse events arrive sequentially, and are aggregated over time. In this work, our goal is to identify whether there are subgroups—defined by any combination of relevant features—that are disproportionately likely to experience harmful interactions with the system. We formalize this problem as a sequential hypothesis test, and identify conditions on reporting behavior that are sufficient for making inferences about disparities in true rates of harm across subgroups. We show that algorithms for sequential hypothesis tests can be applied to this problem with a standard multiple testing correction. We then demonstrate our method on real-world datasets, including mortgage decisions and vaccine side effects; on each, our method (re-)identifies subgroups known to experience disproportionate harm using only a fraction of the data that was initially used to discover them.
@@ 1 Introduction
Figure 1: Overview of incident database framework.
@@
Fairness is an especially salient application for incident reporting systems: while individuals bear the harm, commonly-accepted (and legally-legible) notions of fairness are understood at an aggregate level. In fact, existing examples of (algorithmic) discrimination lawsuits (e.g., Gilbert (2023) in hiring, or Pazanowski (2024) in housing) are often structured as class actions, even as they are initiated by individuals based on their personal experiences. Crucially, individuals themselves may not know whether their experience with the system was inherently problematic, and deserving of redress, until it is placed in context with the experiences of others. On the other hand, while existing incident databases do not typically analyze reporting behavior, it may be necessary to consider reporting more carefully in order for incident databases to be useful for fairness auditing in more general settings, such as for algorithms that make allocation decisions.
In this paper, we consider what a realistic approach to assessing fairness claims from an incident database might look like in practice. We are primarily interested in designing a framework for the general public to report and contest large-scale harms by leveraging reports of individual experience to inform collective evidence of discrimination. To this end, we propose incident databases, which allow individuals to submit reports of negative interactions, as a new mechanism for post-deployment fairness auditing. In particular, we identify conditions on reporting behavior and show how they can be used to to make inferences about rates of true harm in Section 3. Our formalization of the problem allows us to leverage known approaches to sequential hypothesis testing. In Section 4 we show how to instantiate two reasonable algorithms for our test and provide theoretical guarantees for each. Finally, in Section 5, we illustrate the usefulness of our approach using real-world datasets, for applications with known disparity in per-subgroup rates of harm. On both real vaccine incident reports and on mortgage allocation decisions, our algorithm correctly identifies groups that disproportionately experience harm—and does so using a comparatively small number of reports.
@@ 1.1 Related work & application context
The incident database problem is at the intersection of various challenges addressed in fairness and statistics.
@@ 2 Model, Notation, and Preliminaries
The goal of constructing an incident database is to determine whether some system that individuals interact with—for example, an (algorithmic) loan decision system, or a medical treatment—results in disproportionate harm to some meaningful subgroups. For the incident database associated with a particular system, we will use $Y\in\{0,1\}$ as an indicator variable that denotes the undesirable event corresponding to that system. For example, in loan decisions, this could correspond to the event that a highly-qualified individual was denied a loan; in the medical setting, this may be an adverse physical side effect due to the treatment.
@@
Each report $X_{t}$ is received sequentially, and assumed to be sampled i.i.d. from some underlying reporting distribution. [^1] Note that the events $\{X_{t}\in G\}$ and $\{X_{t}\in G^{\prime}\}$ are correlated for any $G,G^{\prime}\in{\mathcal{G}}$, i.e. the independence does not hold across groups. The key point in our case is independence across time. Given a group $G$, we denote its corresponding mean among reports $\Pr[X_{t}\in G\mid R_{t}=1]$ as $\mu_{G}$. We will sometimes refer to $\{\mu_{G}\}_{G\in{\mathcal{G}}}$ as (reporting) preponderances, as they represent the proportion of reports that each $G$ comprises. A central claim of this paper is that comparing $\mu_{G}$ to ${\mu_{G}^{0}}$—i.e., the extent to which group $G$ is (over)represented within the reporting database—can be a useful signal for $\Pr[Y\mid G]$ in a wide class of applications. [^2] Because we allow groups to overlap, we cannot enforce $\sum_{G}{\mu_{G}^{0}}=1$ or $\sum_{G}\mu_{G}=1$. The i.i.d. model of course simplifies the analysis and exposition, but itself is not intrinsic to modeling the incident reporting problem as a sequential hypothesis test. As we discuss in Appendix B, the explicit i.i.d. assumption can be relaxed; more generally, any probabilistic model for sequential testing can be adapted to incident reporting.
@@ 5.1 Myocarditis from COVID-19 vaccines
The Vaccine Adverse Event Reporting System (VAERS) is a national adverse event incident database for U.S.-licensed vaccines, co-managed by the Centers for Disease Control and Prevention (CDC) and the U.S. Food and Drug Administration (FDA) (Chen et al., 1994; Shimabukuro et al., 2015). The database is a combination of voluntary reports from patients that have received the vaccine, as well as mandatory reports from vaccine manufacturers and healthcare professionals. For this case study, we filter the set of publicly-available reports from VAERS to reports about the COVID-19 vaccine with a complaint keyword including “myocarditis.” As for how a database administrator would have known to focus on myocarditis a priori, one might imagine, for example, that the series of case studies found in early 2021 raised the alarm that more systematic analysis was warranted for myocarditis in particular.
@@ 6 Discussion
This work is an initial approach to using incident databases for post-deployment auditing; we believe there is a rich range of future work that develops the ideas in this paper, both technically and conceptually.
@@
More conceptually, while the application examples in Section 5 are somewhat stylized, they demonstrate that incident databases can be promising starting points for new types of post-deployment evaluation. For incident databases to be practically useful, there are a plethora of additional considerations to incorporate from a variety of disciplines. For instance, if a reporting system was available, how would individuals engage with them in theory, and in practice? To what extent do, and should, individual incentives affect the database, and how it is designed? How can the result of a test (a null hypothesis rejection) be contextualized by existing and emerging legal frameworks?
@@ Appendix B Further practical considerations
Note that testing against base preponderances of the reference population (i.e., to compare $\mu_{G}$ to ${\mu_{G}^{0}}$) is a new test proposed by this work, and the analysis in Sections 3.1 and 3.2 is specific to this test. Existing approaches to monitoring in incident databases compare to different baselines, most commonly the historical overall incidence rate for the specific symptom, sometimes by subgroup [Shimabukuro et al., 2015, Kulldorff et al., 2011, Oster et al., 2022]. These baselines could, in principle, be plugged into the algorithms in Section 4, but new analysis for (possibly group-varying) reporting rates would be necessary to draw inferences about analogous quantities of interest (e.g., $\mathrm{RR}$ or $\mathrm{IR}$), as current approaches do not generally consider reporting behavior. In contrast, our modeling allows us to identify what quantities may affect the true incidence rate even if they may be unmeasurable.

### v2

@@ Abstract
When an individual reports a negative interaction with some system, how can their personal experience be contextualized within broader patterns of system behavior? We study the reporting database problem, where individual reports of adverse events arrive sequentially, and are aggregated over time. In this work, our goal is to identify whether there are subgroups—defined by any combination of relevant features—that are disproportionately likely to experience harmful interactions with the system. We formalize this problem as a sequential hypothesis test, and identify conditions on reporting behavior that are sufficient for making inferences about disparities in true rates of harm across subgroups. We show that algorithms for sequential hypothesis tests can be applied to this problem with a standard multiple testing correction. We then demonstrate our method on real-world datasets, including mortgage decisions and vaccine side effects; on each, our method (re-)identifies subgroups known to experience disproportionate harm using only a fraction of the data that was initially used to discover them.
@@ 1 Introduction
Figure 1: Overview of reporting database framework.
@@
Fairness is an especially salient application for individual reporting systems: while individuals bear the harm, commonly-accepted (and legally-legible) notions of fairness are understood at an aggregate level. In fact, existing examples of (algorithmic) discrimination lawsuits (e.g., Gilbert (2023) in hiring, or Pazanowski (2024) in housing) are often structured as class actions, even as they are initiated by individuals based on their personal experiences. Crucially, individuals themselves may not know whether their experience with the system was inherently problematic, and deserving of redress, until it is placed in context with the experiences of others. On the other hand, while existing reporting databases do not typically analyze reporting behavior, it may be necessary to consider reporting more carefully in order for reporting databases to be useful for fairness auditing in more general settings, such as for algorithms that make allocation decisions.
In this paper, we consider what a realistic approach to assessing fairness claims from an incident reporting database might look like in practice. We are primarily interested in designing a framework for the general public to report and contest large-scale harms by leveraging reports of individual experience to inform collective evidence of discrimination. To this end, we propose reporting databases, which allow individuals to submit reports of negative interactions, as a new mechanism for post-deployment fairness auditing. In particular, we identify conditions on reporting behavior and show how they can be used to to make inferences about rates of true harm in Section 3. Our formalization of the problem allows us to leverage known approaches to sequential hypothesis testing. In Section 4 we show how to instantiate two reasonable algorithms for our test and provide theoretical guarantees for each. Finally, in Section 5, we illustrate the usefulness of our approach using real-world datasets, for applications with known disparity in per-subgroup rates of harm. On both real vaccine incident reports and on mortgage allocation decisions, our algorithm correctly identifies groups that disproportionately experience harm—and does so using a comparatively small number of reports.
@@ 1.1 Related work & application context
The reporting database problem is at the intersection of various challenges addressed in fairness and statistics.
@@ 2 Model, Notation, and Preliminaries
The goal of constructing a reporting database is to determine whether some system that individuals interact with—for example, an (algorithmic) loan decision system, or a medical treatment—results in disproportionate harm to some meaningful subgroups. For the reporting database associated with a particular system, we will use $Y\in\{0,1\}$ as an indicator variable that denotes the undesirable event corresponding to that system. For example, in loan decisions, this could correspond to the event that a highly-qualified individual was denied a loan; in the medical setting, this may be an adverse physical side effect due to the treatment.
@@
Each report $X_{t}$ is received sequentially, and assumed to be sampled i.i.d. from some underlying reporting distribution. [^1] Note that the events $\{X_{t}\in G\}$ and $\{X_{t}\in G^{\prime}\}$ are correlated for any $G,G^{\prime}\in{\mathcal{G}}$, i.e. the independence does not hold across groups. The key point in our case is independence across time. Given a group $G$, we denote its corresponding mean among reports $\Pr[X_{t}\in G\mid R_{t}=1]$ as $\mu_{G}$. We will sometimes refer to $\{\mu_{G}\}_{G\in{\mathcal{G}}}$ as (reporting) preponderances, as they represent the proportion of reports that each $G$ comprises. A central claim of this paper is that comparing $\mu_{G}$ to ${\mu_{G}^{0}}$—i.e., the extent to which group $G$ is (over)represented within the reporting database—can be a useful signal for $\Pr[Y\mid G]$ in a wide class of applications. [^2] Because we allow groups to overlap, we cannot enforce $\sum_{G}{\mu_{G}^{0}}=1$ or $\sum_{G}\mu_{G}=1$. The i.i.d. model of course simplifies the analysis and exposition, but itself is not intrinsic to modeling the reporting database problem as a sequential hypothesis test. As we discuss in Appendix B, the explicit i.i.d. assumption can be relaxed; more generally, any probabilistic model for sequential testing can be adapted to reporting databases.
@@ 5.1 Myocarditis from COVID-19 vaccines
The Vaccine Adverse Event Reporting System (VAERS) is a national adverse event incident reporting database for U.S.-licensed vaccines, co-managed by the Centers for Disease Control and Prevention (CDC) and the U.S. Food and Drug Administration (FDA) (Chen et al., 1994; Shimabukuro et al., 2015). The database is a combination of voluntary reports from patients that have received the vaccine, as well as mandatory reports from vaccine manufacturers and healthcare professionals. For this case study, we filter the set of publicly-available reports from VAERS to reports about the COVID-19 vaccine with a complaint keyword including “myocarditis.” As for how a database administrator would have known to focus on myocarditis a priori, one might imagine, for example, that the series of case studies found in early 2021 raised the alarm that more systematic analysis was warranted for myocarditis in particular.
@@ 6 Discussion
This work is an initial approach to using reporting databases for post-deployment auditing; we believe there is a rich range of future work that develops the ideas in this paper, both technically and conceptually.
@@
More conceptually, while the application examples in Section 5 are somewhat stylized, they demonstrate that reporting databases can be promising starting points for new types of post-deployment evaluation. For reporting databases to be practically useful, there are a plethora of additional considerations to incorporate from a variety of disciplines. For instance, if a reporting system was available, how would individuals engage with them in theory, and in practice? To what extent do, and should, individual incentives affect the database, and how it is designed? How can the result of a test (a null hypothesis rejection) be contextualized by existing and emerging legal frameworks?
@@ Appendix B Further practical considerations
Note that testing against base preponderances of the reference population (i.e., to compare $\mu_{G}$ to ${\mu_{G}^{0}}$) is a new test proposed by this work, and the analysis in Sections 3.1 and 3.2 is specific to this test. Existing approaches to monitoring in reporting databases compare to different baselines, most commonly the historical overall incidence rate for the specific symptom, sometimes by subgroup [Shimabukuro et al., 2015, Kulldorff et al., 2011, Oster et al., 2022]. These baselines could, in principle, be plugged into the algorithms in Section 4, but new analysis for (possibly group-varying) reporting rates would be necessary to draw inferences about analogous quantities of interest (e.g., $\mathrm{RR}$ or $\mathrm{IR}$), as current approaches do not generally consider reporting behavior. In contrast, our modeling allows us to identify what quantities may affect the true incidence rate even if they may be unmeasurable.